OpenStudy (linn99123):

NEED HELP FAST WILL FAN AND GIVE MEDAL A 12-foot ladder is leaning against a wall. The bottom of the ladder is 5 feet away from the bottom of the wall. Approximately how high up the wall does the top of the ladder reach? A. 2.4 feet B. 7.0 feet C. 13.0 feet D. 10.9 feet

OpenStudy (davidusa):

10.9

OpenStudy (davidusa):

ok?

OpenStudy (linn99123):

u shure?

OpenStudy (davidusa):

yes

OpenStudy (davidusa):

100 percent sure

OpenStudy (anonymous):

10.9 looks to be the best choice

OpenStudy (johnweldon1993):

 Remember Pythagorean theorem c^2 = a^2 + b^2 You have C and you have A... 5^2 + b^2 = 12^2 25 + b^2 = 144 b^2 = 119 b = ?

OpenStudy (davidusa):

yep i got it right

random231 (random231):

yup

OpenStudy (calculusxy):

So you wanna find the squre root of 119.